OpenStudy (anonymous):

The formula p = 2ℓ + 2w represents the perimeter of a rectangle. In this formula, ℓ is the length of the rectangle and w is the width. Solve the formula for ℓ. Find the length when the width is 4 meters and the perimeter is 36 meters.

OpenStudy (anonymous):

p = 2l + 2w making l the subject of formula, l = p-2w/2 when p = 36 and w = 4 l =36 - 2(4)/2 l = 14 meters

OpenStudy (anonymous):

@kaytiemaguire